# Periodic orbits and gravitational wave radiation in short hair black hole spacetimes for an extreme mass ratio system

### v2

@@ Abstract
For a short hair black hole(BH) which circumvents the “no-short-hair” theorem, it manifests intense hair behavior in the vicinity of the event horizon, accompanied by remarkable quantum effects. These effects may carry important information about the internal structure and dynamical evolution of BHs, thereby providing a new perspective on the problem of black hole information loss. Therefore, in this paper, we analyze the influence of the hair parameter $Q_{m}$ and the structural parameter $k$ of the short hair BH in an extreme mass ratio(EMR) system on the periodic orbits of particles and gravitational wave radiation. The results show that as $Q_{m}$ increases, the radius and angular momentum of the bound orbit decrease, and the $E-L$ space shifts to the left. An increase in $k$ weakens this trend. When $k$ takes a larger value, the short hair BH and the Schwarzschild BH tend to be degenerate. Compared with the Schwarzschild BH, the bound orbit energy and angular momentum of the short hair black hole are reduced. Under the conditions of higher $Q_{m}$ and lower $k$, the gravitational wave period is suppressed and the signal amplitude is significantly increased. These results provide new observational means for distinguishing short hair BH from classical BH and offer new insights for verifying the no-hair theorem and understanding the physical behavior near the event horizon.
@@ I Introduction
General relativity(GR) successfully predicts the existence of BHs and is supported by a wealth of astrophysical observations. For instance, the gravitational wave signals produced by the merger of binary BHs Abbott et al. (2016a, b, 2019, 2020), as well as the capture of BH shadow images and the observation of the supermassive BH at the center of the Milky Way Akiyama et al. (2019, 2022); Genzel et al. (2010). These discoveries strongly confirm the success of GR. However, within the framework of classical GR, the inevitability of spacetime singularities Penrose (1965) and the problem of information loss caused by the event horizon Hawking (2005) reveal the limitations of GR when dealing with extreme gravitational fields. These limitations prompt researchers to explore possible modifications or develop new gravitational theories, such as loop quantum gravity Ashtekar et al. (2006a, b, 2003); Singh and Toporensky (2004); Wilson-Ewing (2017); Lewandowski et al. (2023), string theory Witten (1995); Horowitz and Polchinski (1997); Brandenberger and Vafa (1989), and theories of the coupling of gravity and matter fields Bizon (1990); Brown and Husain (1997); Bakopoulos and Nakas (2023). These theories provide new perspectives for understanding phenomena in extreme gravitational fields. However, verifying the rationality of these theories requires a suitable experimental platform, and the extreme gravitational field environment around BHs is an ideal natural experimental site for testing these new theories. The extreme gravitational field near BHs can not only generate gravitational waves and BH shadows but also be accompanied by many complex astronomical phenomena, including gravitational lensing effects Zhao et al. (2024a, b); Tsukamoto (2017), photon rings Yang et al. (2024a), and bound orbits. These phenomena provide extremely rich information for studying the spacetime structure near BHs and provide important opportunities for testing gravitational theories in strong gravitational fields.The motion of massless photons along null geodesics determines the shape of the black hole shadow and is closely related to the gravitational lensing effect. The timelike geodesics of massive particles describe the orbital evolution of objects in strong gravitational fields, and are particularly important for characterizing the gravitational wave radiation in binary systems and EMR systems. Therefore, in-depth study of geodesics not only helps to test GR and its modified theories but also helps to further understand the behavior near BHs, especially near the event horizon.
@@
As is well known, the bound orbits of timelike geodesic particles can be divided into precessing orbits and periodic orbits. The early precession of Mercury’s perihelion Will (1982) provided one of the early evidences for the success of GR. Its precession results from the weak effect of the solar gravitational field. In strong gravitational fields, the precession phenomenon of stars orbiting supermassive BHs has been widely studied, especially the stellar orbits around Sgr A*. The references are Hees et al. (2017); Will (2008); Iorio (2011); Grould et al. (2017); Amorim et al. (2019); Abuter et al. (2020). Among them, the observation of the precession of the star S2 around the massive Sgr A* BH Abuter et al. (2020) has been widely applied in testing other theories and constraining BH parameters Hees et al. (2017); Qi et al. (2024); Della Monica and de Martino (2023); Li et al. (2024).In contrast, periodic orbits are more complex. The motion characteristics of these orbits often exhibit zoom-whirl behavior. Therefore, they can encode the orbital information around BHs through the orbits Levin and Perez-Giz (2008). These periodic orbits are often classified by the orbital classification method in reference Levin and Perez-Giz (2008). In their literature, it is pointed out that periodic orbits can often use rational or irrational numbers $q$ to encode their orbits. Here, $q$ is related to the angular frequency and the radial frequency. When q is an irrational number, the bound orbit at this time shows non-periodic orbits and precession behavior. When $q$ is a rational number, the orbit at this time shows periodic orbits. The orbit corresponding to an irrational number can always be well approximated by a periodic orbit Levin and Perez-Giz (2008). This also leads people to generally believe that periodic orbits can provide more information about the strong field region than precession. Therefore, a large number of studies on periodic orbits using classification methods have emerged. For example, studies on periodic orbits around Schwarzschild BH, Reissner-Nordström BH, and Kerr BH Lim and Yeo (2024); Levin and Perez-Giz (2008); Misra and Levin (2010); Levin (2009); Grossman et al. (2012, 2013). In addition, studies on periodic orbits in other spacetime backgrounds are also gradually enriched Lin and Deng (2023); Deng (2020); Zhou and Xie (2020); Wei et al. (2019); Mummery and Balbus (2022); Babar et al. (2017).
@@
Furthermore, for a short hair BH that violates the no-hair theorem and bypasses the ”no-short-hair” theorem Brown and Husain (1997), it is characterized by generating arbitrarily short hair near the event horizon by considering the coupling of anisotropic matter with Einstein’s gravity. These hairs are usually accompanied by quantum effects and may carry information about the internal structure and evolution of the BH Brown and Husain (1997). Especially near the event horizon, quantum fluctuations and gravitational perturbations may have a significant impact on the local spacetime. Therefore, by studying timelike geodesics, periodic orbits and their gravitational wave radiation, the properties of these short hairs and their impact on the spacetime geometry near the BH’s horizon can be deeply understood. Such gravitational wave signals not only provide a new observational window for BH dynamics but also may reveal the internal information of the BH and its evolution process, thereby providing new insights for the research of the BH information paradox and quantum gravity.
@@ II Short Hair Black Hole and Timelike Geodesics
The no-hair theorem states that a BH can be described by just three parameters: mass $M$, angular momentum $J$, and electric charge $Q$ Ruffini and Wheeler (1971); Hawking (1972); Israel (1967). However, when the exterior of a BH is filled with various matter fields, the no-hair theorem will be violated. The first BH solution that violates the no-hair theorem was calculated by Piotr Bizon. He obtained a “charged” BH solution through numerical analysis methods Bizon (1990). The emergence of this BH solution that violates the no-hair theorem has attracted extensive attention in the academic community and prompted researchers to actively explore more BH solutions with hair. In this regard, Ovalle and Contreras et al. successfully constructed static spherically symmetric and axially symmetric hairy BH solutions through the gravitational decoupling method Ovalle et al. (2021); Contreras et al. (2021). Further research has obtained more BH solutions with hair by introducing different types of matter fields. These achievements are reported in relevant literatures respectively Caramês (2023); Bakopoulos and Nakas (2023); Zhang (2022); Brihaye and Ducobu (2016); Lavrelashvili and Maison (1993); Hod (2011); Tahamtan (2020); Creminelli et al. (2020); Herdeiro and Radu (2020); Clough et al. (2019); Brown and Husain (1997). Among them, the authors of reference Brown and Husain (1997) obtained a class of static spherically symmetric BH solutions with short hair by considering the coupling of Einstein gravity and anisotropic matter fields. The hair of this kind of solution only has a significant influence near the event horizon and does not violate the ”no-short-hair” theorem. The detailed reasons will be discussed later.
@@
$\begin{cases}g_{rr}=0\\ \frac{dg_{rr}}{dr}|_{r=r_{h}}=0\end{cases}.$ (5)
@@
Here, we stipulate that the hairy parameter $Q_{m}=Q_{c}$ for the extreme black hole case. Therefore, the value range of the hairy parameter is $0\leq Q_{m}\leq Q_{c}$. As shown in Fig. 1, the change in the number of event horizons of the short hair BH when the hair parameter $Q_{m}$ takes different values. Obviously, when the hair parameter is $Q_{c}$, the short hair BH at this time is an extreme BH (the Cauchy horizon and the event horizon are degenerate). When the hair parameter $Q_{m}=0$, the BH at this time degenerates into a Schwarzschild BH. The black dashed line in the figure represents the Schwarzschild BH case. It is easy to obtain from the figure that the existence of hair makes the event horizon radius smaller than the event horizon radius of the Schwarzschild BH. When the hair parameter takes a value of $Q_{m}<Q_{c}$, there are two horizons for the short hair BH at this time. When the hair parameter takes a value of $Q_{m}>Q_{c}$, there is no BH in the spacetime at this time, but a naked singularity appears. In this article, we focus on the case where the short hair BH always exists. Therefore, the value range of the hairy parameter is $0\leq Q_{m}\leq Q_{c}$.
@@ III Bound Orbits of Test Particles in Short Hair Black Hole Spacetimes
In the context of a short hair BH spacetime, periodic orbits in an EMR system exhibit a range of unique physical characteristics. The orbital structure and evolutionary behavior of stellar-mass compact objects, influenced by the strong gravitational field of the primary BH, differ significantly from those in conventional mass ratio systems. Notably, in the modified spacetime background of a short hair BH, these orbits may display periodic behaviors and orbital evolutions that are markedly different from those predicted by classical BH models, thereby having significant implications for the overall dynamics of the system and the characteristics of gravitational wave radiation. In this section, we will explore the properties of these periodic orbits. We adopt the classification scheme from the literature Levin and Perez-Giz (2008) to describe these orbits $(z,w,v)$ using a unique irrational or rational number $q$. The expression is given by:
@@
When $q$ is an irrational number, the orbit of a neutral particle at this time exhibits precession phenomena, similar to the perihelion precession of Mercury. When $q$ is a rational number, the motion of the particle at this time shows periodic characteristics, namely a periodic orbit. As pointed out in reference Levin and Perez-Giz (2008), any general orbit can be approximated by a periodic orbit. Therefore, the study of periodic orbits is helpful for understanding the properties of the BH background and has potential application value in understanding the gravitational radiation mechanism around BHs. In expression (15), $w$ represents the whirl number of the orbit, $v$ denotes the vertex number of the orbit, and $z$ signifies the zoom number (or blade number). It is worth noting that these three elements are all integers. $\Delta\phi$ represents the angular change within each period. Its expression is
@@
Fig. 7 shows the variation of the rational number $q$ with $E$ or $L$. In the first row of figures, since the orbital angular momentum range is $L_{ISCO}<L<L_{MBO}$, the angular momentum $L$ can be expressed as $L=L_{ISCO}+\varepsilon(L_{MBO}-L_{ISCO})$. Without loss of generality, we fix $\varepsilon=\frac{1}{2}$ to plot the curve of $q$ varying with E (see Fig. 7). Apparently, within the energy range of $E_{ISCO}<E<E_{MBO}$, the rational or irrational number $q$ under the background of three types of short hair BHs starts to increase slowly with the increase of energy and rises sharply to divergence near the maximum energy value. The existence of hair causes the curve to shift to the left and diverge at the maximum value of its corresponding energy, and the greater the hair strength, the faster it will reach the divergence point. This is consistent with the transformation situation of the $E-L$ space shown in Fig. 5. That is to say, the existence of hair makes it diverge earlier compared with the Schwarzschild BH case, and the greater the hair strength, the faster the divergence. The blue curve in the figure represents the situation of the Schwarzschild BH. It is worth mentioning that we find that when the parameter $k$ increases, the behavior of the hair parameter on $q$ will approach the situation of the Schwarzschild BH and eventually degenerate when $k$ takes a larger value. In the second row of figures, we fix the energy $E=0.96$ to plot the variation of $q$ with the angular momentum $L$. Apparently, as the angular momentum decreases, $q$ first increases slowly. When the angular momentum decreases to near its minimum value, $q$ increases sharply and eventually diverges. The existence of hair parameters makes it diverge earlier, and the stronger the intensity, the more obvious it is. As the value of $k$ increases, these trends will approach the situation of the Schwarzschild BH and eventually degenerate (the blue curve in the figure represents the situation of the Schwarzschild BH).
@@
$h_{ij}=\frac{4\mu M}{D_{L}}\left(v_{i}v_{j}-\frac{m}{r}n_{i}n_{j}\right),$ (18)
where $M$ represents the mass of the supermassive BH (here it is the mass of the short hair BH), $m$ is the mass of the small-mass celestial body, and $D_{L}$ is the luminosity distance of the EMR system. $\mu=\frac{Mm}{\left(M+m\right)^{2}}$ is the symmetric mass ratio of the system. $v_{i}$ represents the velocity vector component, and $n_{i}$ is the unit vector pointing in the radial direction.
Next, we can project this gravitational wave onto the detector-adapted coordinates Poisson and Will (2014). The coordinates are as follows:
@@
Under such a coordinate basis, the polarization components $h_{+}$ and $h_{\times}$ can be expressed as Poisson and Will (2014)
@@
where $\iota$ is the inclination angle of the test particle’s orbit with respect to the line of sight, and $\phi$ and $\zeta$ represent the phase angle and the latitude angle, respectively, which are related to the orbital phase of the gravitational wave source and the direction of the detector.
@@
In general, the hair parameter $Q_{m}$ and parameter $k$ exert a remarkable influence on the gravitational wave signals radiated by the periodic orbits. Specifically, a larger value of the hair parameter $Q_{m}$ and a smaller value of parameter $k$ will suppress the period of the gravitational wave signal, simultaneously increase the amplitude, and alter its phase. This implies that by identifying these differences, future gravitational wave detections can disclose the features of the zoom and whirl behaviors of the orbit. These features not only offer crucial evidence for identifying the properties of the short hair BH, distinguishing it from the Schwarzschild BH, but also open a potential observational window for verifying the no-hair theorem or the ”no-short-hair” theorem. It is worth noting that the adiabatic approximation method has been adopted in this study to calculate the gravitational wave waveform. Although this method cannot capture all the details of the gravitational wave, within a single period, it can effectively reflect the key physical features of the gravitational wave signal, especially in the zoom and whirl behaviors. Therefore, within the scope of the current research, the adiabatic approximation is reasonable and sufficient, especially when analyzing the hair property of the short hair BH and its difference from the Schwarzschild BH. Such a difference is adequate. However, the adiabatic approximation ignores the non-adiabatic effects and high-order multipole moment contributions in the orbital evolution. In order to further study the impact of these effects on the gravitational wave waveform and more accurately characterize the unique features of the short hair BH, a more accurate waveform model needs to be introduced in future work. This will not only help deepen the understanding of the radiation of periodic orbits but also provide new observational opportunities for distinguishing the short hair BH from the classical BH. And these contents will be further explored in future research.

### v3

@@ Abstract
For a short hair black hole(BH) that bypasses the “no-short-hair” theorem, its hair parameters have a significant impact near the event horizon and may thus carry the internal information of the BH. This paper studies the influence of the hair parameter $Q_{m}$ and the structural parameter $k$ of the short hair BH on the periodic orbits of particles and gravitational wave radiation in the extreme mass ratio(EMR) system. The research results show that $Q_{m}$ has a significant impact on the radius of the orbit, angular momentum and the $E-L$ space. A larger value of $k$ weakens this effect and makes it degenerate with the Schwarzschild BH. In addition, an increase in $Q_{m}$ or a decrease in $k$ will lead to a reduction in the period of gravitational waves and a significant enhancement of the signal amplitude. These results provide new observational means for distinguishing short hair BH from classical BH, and also offer new insights for verifying the no-hair theorem and understanding the physical behaviors near the event horizon.
@@ I Introduction
General relativity(GR) successfully predicts the existence of BHs and is supported by a wealth of astrophysical observations. For instance, the gravitational wave signals produced by the merger of binary BHs Abbott et al. (2016a, b, 2019, 2020), as well as the capture of BH shadow images and the observation of the supermassive BH at the center of the Milky Way Akiyama et al. (2019, 2022); Genzel et al. (2010). These discoveries strongly confirm the success of GR. However, within the framework of classical GR, the inevitability of spacetime singularities Penrose (1965) and the problem of information loss caused by the event horizon Hawking (2005) reveal the limitations of GR when dealing with extreme gravitational fields. These limitations prompt researchers to explore possible modifications or develop new gravitational theories, such as loop quantum gravity (see, e.g.,Ashtekar et al. (2006a, b, 2003); Singh and Toporensky (2004); Wilson-Ewing (2017); Lewandowski et al. (2023)), string theory (see, e.g., Witten (1995); Horowitz and Polchinski (1997); Brandenberger and Vafa (1989)), and theories of the coupling of gravity and matter fields (see, e.g.,Bizon (1990); Brown and Husain (1997); Bakopoulos and Nakas (2023)). These theories provide new perspectives for understanding phenomena in extreme gravitational fields. However, verifying the rationality of these theories requires a suitable experimental platform, and the extreme gravitational field environment around BHs is an ideal natural experimental site for testing these new theories. The extreme gravitational field near BHs can not only generate gravitational waves and BH shadows but also be accompanied by many complex astronomical phenomena, including gravitational lensing effects (see, e.g.,Zhao et al. (2024a, b); Tsukamoto (2017), photon rings Yang et al. (2024a)), and bound orbits. These phenomena provide extremely rich information for studying the spacetime structure near BHs and provide important opportunities for testing gravitational theories in strong gravitational fields.The motion of massless photons along null geodesics determines the shape of the black hole shadow and is closely related to the gravitational lensing effect. The timelike geodesics of massive particles describe the orbital evolution of objects in strong gravitational fields, and are particularly important for characterizing the gravitational wave radiation in binary systems and EMR systems. Therefore, in-depth study of geodesics not only helps to test GR and its modified theories but also helps to further understand the behavior near BHs, especially near the event horizon.
@@
As is well known, the bound orbits of timelike geodesic particles can be divided into precessing orbits and periodic orbits. The early precession of Mercury’s perihelion Will (1982) provided one of the early evidences for the success of GR. Its precession results from the weak effect of the solar gravitational field. In strong gravitational fields, the precession phenomenon of stars orbiting supermassive BHs has been widely studied, especially the stellar orbits around Sgr A*. The references are (see, e.g.,Hees et al. (2017); Will (2008); Iorio (2011); Grould et al. (2017); Amorim et al. (2019); Abuter et al. (2020)). Among them, the observation of the precession of the star S2 around the massive Sgr A* BH Abuter et al. (2020) has been widely applied in testing other theories and constraining BH parameters Hees et al. (2017); Qi et al. (2024); Della Monica and de Martino (2023); Li et al. (2024).In contrast, periodic orbits are more complex. The motion characteristics of these orbits often exhibit zoom-whirl behavior. Therefore, they can encode the orbital information around BHs through the orbits Levin and Perez-Giz (2008). These periodic orbits are often classified by the orbital classification method in reference Levin and Perez-Giz (2008). In their literature, it is pointed out that periodic orbits can often use rational or irrational numbers $q$ to encode their orbits. Here, $q$ is related to the angular frequency and the radial frequency. When q is an irrational number, the bound orbit at this time shows non-periodic orbits and precession behavior. When $q$ is a rational number, the orbit at this time shows periodic orbits. The orbit corresponding to an irrational number can always be well approximated by a periodic orbit Levin and Perez-Giz (2008). This also leads people to generally believe that periodic orbits can provide more information about the strong field region than precession. Therefore, a large number of studies on periodic orbits using classification methods have emerged. For example, studies on periodic orbits around Schwarzschild BH, Reissner-Nordström BH, and Kerr BH Lim and Yeo (2024); Levin and Perez-Giz (2008); Misra and Levin (2010); Levin (2009); Grossman et al. (2012, 2013). In addition, studies on periodic orbits in other spacetime backgrounds are also gradually enriched (see, e.g.,Lin and Deng (2023); Deng (2020); Zhou and Xie (2020); Wei et al. (2019); Mummery and Balbus (2022); Babar et al. (2017)).
@@
Furthermore, for a short hair BH that violates the no-hair theorem and circumvents the “no-short-hair” theorem Brown and Husain (1997), its main feature lies in generating arbitrarily short ”hairs” near the event horizon by taking into account the coupling between anisotropic matter and the Einstein gravitational field. These hairs are usually accompanied by quantum effects and may carry important information about the internal structure and evolution of BHs Brown and Husain (1997). Especially near the event horizon, quantum fluctuations and gravitational perturbations may significantly affect the local spacetime structure. Therefore, studying timelike geodesics, periodic orbits and their gravitational wave radiation characteristics in this spacetime background will help to gain a deeper understanding of the properties of the short hair BH.
This paper mainly focuses on comparing the dynamical behaviors of classical Schwarzschild BHs and short hair BH in the EMR system. By thoroughly analyzing the differences between these two types of BH in terms of timelike geodesics, periodic orbits and the characteristics of gravitational wave radiation, it aims to reveal the unique manifestations of short hair BHs under extreme gravitational fields. Such a comparison not only helps to evaluate the applicability of the short hair BH model in explaining observational phenomena but also provides a new theoretical basis for testing the modified theories of general relativity. Specifically, we explore the characteristics of gravitational wave signals of the short hair BH in the EMR system and conduct a comparative analysis with the signals of the classical Schwarzschild BH to investigate the influence of hair parameter on these signals. Through this comparative study, it is expected to provide some theoretical references for future astronomical observations and promote a deeper understanding of the extreme gravitational field environment around BHs. Such gravitational wave signals not only offer a new observational window for BH dynamics but may also reveal the internal information of BHs and their evolution processes, thereby providing a new perspective for resolving the BH information paradox and advancing the development of the quantum gravity theory.
@@ II Short Hair Black Hole and Timelike Geodesics
The no-hair theorem states that a BH can be described by just three parameters: mass $M$, angular momentum $J$, and electric charge $Q$ Ruffini and Wheeler (1971); Hawking (1972); Israel (1967). However, when the exterior of a BH is filled with various matter fields, the no-hair theorem will be violated. The first BH solution that violates the no-hair theorem was calculated by Piotr Bizon. He obtained a “charged” BH solution through numerical analysis methods Bizon (1990). The emergence of this BH solution that violates the no-hair theorem has attracted extensive attention in the academic community and prompted researchers to actively explore more BH solutions with hair. In this regard, Ovalle and Contreras et al. successfully constructed static spherically symmetric and axially symmetric hairy BH solutions through the gravitational decoupling method Ovalle et al. (2021); Contreras et al. (2021). Further research has obtained more BH solutions with hair by introducing different types of matter fields. These achievements are reported in relevant literatures respectively (see, e.g., Caramês (2023); Bakopoulos and Nakas (2023); Zhang (2022); Brihaye and Ducobu (2016); Lavrelashvili and Maison (1993); Hod (2011); Tahamtan (2020); Creminelli et al. (2020); Herdeiro and Radu (2020); Clough et al. (2019); Brown and Husain (1997)). Among them, the authors of reference Brown and Husain (1997) obtained a class of static spherically symmetric BH solutions with short hair by considering the coupling of Einstein gravity and anisotropic matter fields. The hair of this kind of solution only has a significant influence near the event horizon and does not violate the ”no-short-hair” theorem. The detailed reasons will be discussed later.
@@
$\begin{cases}g^{rr}=0\\ \frac{dg^{rr}}{dr}|_{r=r_{h}}=0\end{cases}.$ (5)
@@
Here, we stipulate that the hair parameter $Q_{m}$ of the short hair BH reaches a critical value $Q_{c}$, that is, $Q_{m}=Q_{c}$. This condition corresponds to the situation of an extremal BH. Therefore, the value range of the hairy parameter is $0\leq Q_{m}\leq Q_{c}$. As shown in Fig. 1, the change in the number of event horizons of the short hair BH when the hair parameter $Q_{m}$ takes different values. Obviously, when the hair parameter is $Q_{c}$, the short hair BH at this time is an extreme BH (the Cauchy horizon and the event horizon are degenerate). When the hair parameter $Q_{m}=0$, the BH at this time degenerates into a Schwarzschild BH. The black dashed line in the figure represents the Schwarzschild BH case. It is easy to obtain from the figure that the existence of hair makes the event horizon radius smaller than the event horizon radius of the Schwarzschild BH. When the hair parameter takes a value of $Q_{m}<Q_{c}$, there are two horizons for the short hair BH at this time. When the hair parameter takes a value of $Q_{m}>Q_{c}$, there is no BH in the spacetime at this time, but a naked singularity appears. In this article, we focus on the case where the short hair BH always exists. Therefore, the value range of the hairy parameter is $0\leq Q_{m}\leq Q_{c}$.
@@ III Bound Orbits of Test Particles in Short Hair Black Hole Spacetimes
In the context of a short hair BH spacetime, periodic orbits in an EMR system exhibit a range of unique physical characteristics. The orbital structure and evolutionary behavior of stellar-mass compact objects, influenced by the strong gravitational field of the primary BH, differ significantly from those in conventional mass ratio systems. Notably, in the modified spacetime background of a short hair BH, these orbits may display periodic behaviors and orbital evolutions that are markedly different from those predicted by classical BH models, thereby having significant implications for the overall dynamics of the system and the characteristics of gravitational wave radiation. In this section, we will explore the properties of these periodic orbits. We adopt the classification scheme from the literature Levin and Perez-Giz (2008) to describe these orbits $(z,w,v)$ using a unique irrational or rational number $q$. The expression is given by
@@
In expression (15), $w$ represents the whirl number of the orbit, $v$ denotes the vertex number of the orbit, and $z$ signifies the zoom number (or blade number). It is worth noting that these three elements are all integers. When $q$ is an irrational number, the orbit of a neutral particle at this time exhibits precession phenomena, similar to the perihelion precession of Mercury. When $q$ is a rational number, the motion of the particle at this time shows periodic characteristics, namely a periodic orbit. As pointed out in reference Levin and Perez-Giz (2008), any general orbit can be approximated by a periodic orbit. Therefore, the study of periodic orbits is helpful for understanding the properties of the BH background and has potential application value in understanding the gravitational radiation mechanism around BHs. $\Delta\phi$ represents the angular change within each period. Its expression is
@@
Here we use numerical methods to plot the variation of the rational number $q$ with respect to $E$ or $L$. As shown in Fig. 7. In the first row of figures, since the orbital angular momentum range is $L_{ISCO}<L<L_{MBO}$, the angular momentum $L$ can be expressed as $L=L_{ISCO}+\varepsilon(L_{MBO}-L_{ISCO})$. Without loss of generality, we fix $\varepsilon=\frac{1}{2}$ to plot the curve of $q$ varying with E (see Fig. 7). Apparently, within the energy range of $E_{ISCO}<E<E_{MBO}$, the rational or irrational number $q$ under the background of three types of short hair BHs starts to increase slowly with the increase of energy and rises sharply to divergence near the maximum energy value. The existence of hair causes the curve to shift to the left and diverge at the maximum value of its corresponding energy, and the greater the hair strength, the faster it will reach the divergence point. This is consistent with the transformation situation of the $E-L$ space shown in Fig. 5. That is to say, the existence of hair makes it diverge earlier compared with the Schwarzschild BH case, and the greater the hair strength, the faster the divergence. The blue curve in the figure represents the situation of the Schwarzschild BH. It is worth mentioning that we find that when the parameter $k$ increases, the behavior of the hair parameter on $q$ will approach the situation of the Schwarzschild BH and eventually degenerate when $k$ takes a larger value. In the second row of figures, we fix the energy $E=0.96$ to plot the variation of $q$ with the angular momentum $L$. Apparently, as the angular momentum decreases, $q$ first increases slowly. When the angular momentum decreases to near its minimum value, $q$ increases sharply and eventually diverges. The existence of hair parameters makes it diverge earlier, and the stronger the intensity, the more obvious it is. As the value of $k$ increases, these trends will approach the situation of the Schwarzschild BH and eventually degenerate (the blue curve in the figure represents the situation of the Schwarzschild BH).
@@
$h_{ij}=\frac{4\mu M}{D_{L}}\left(v_{i}v_{j}-\frac{m}{r}n_{i}n_{j}\right).$ (18)
In Eq. (18), $M$ represents the mass of the supermassive BH (here, the mass of the short hair BH), and $m$ is the mass of the small celestial body. Considering that we are studying an EMR system, thus $m\ll M$. $D_{L}$ is the luminosity distance of the EMR system. $\mu=\frac{Mm}{(M+m)^{2}}$ is the symmetric mass ratio of the system. $v_{i}$ represents the component of the velocity vector of the small celestial body and describes the motion state of the small celestial body on its orbit; $n_{i}$ is the unit vector pointing in the radial direction, that is, the direction vector pointing from the BH to the small celestial body, which is obtained after normalization processing.
The quadrupole moment formula (18) describes the spacetime perturbation of the gravitational wave generated by the motion of a small celestial body in the gravitational field of a supermassive BH. This waveform provides the basic distribution characteristics of the gravitational wave in space. However, in order to match it with the actual detector, we need to match the calculated gravitational wave signal with the actual detector, which requires projecting the gravitational wave onto the coordinate system suitable for the detector. As shown in the literature Poisson and Will (2014), its coordinate basis can be written as
@@
In the above coordinate basis, $\zeta$ is the longitude of periastron, $\iota$ is the inclination angle of the test particle orbit relative to the observation direction, and $e_{X}$, $e_{Y}$, $e_{Z}$ define the orthogonal coordinate basis suitable for the detector and are used to decompose the gravitational wave signal into different polarization components. Under such a coordinate basis, the polarization components $h_{+}$ and $h_{\times}$ can be expressed as Poisson and Will (2014)
@@
here, $\phi$ represents the phase angle.
@@
In general, the hair parameter $Q_{m}$ and parameter $k$ exert a remarkable influence on the gravitational wave signals radiated by the periodic orbits. Specifically, a larger value of the hair parameter $Q_{m}$ and a smaller value of parameter $k$ will suppress the period of the gravitational wave signal, simultaneously increase the amplitude, and alter its phase. This implies that by identifying these differences, future gravitational wave detections can disclose the features of the zoom and whirl behaviors of the orbit. These features not only offer crucial evidence for identifying the properties of the short hair BH, distinguishing it from the Schwarzschild BH, but also open a potential observational window for verifying the no-hair theorem or the ”no-short-hair” theorem.
It is worth noting that this paper employs the adiabatic approximation method to calculate the gravitational wave waveform. Although this method is unable to fully capture all the details of the gravitational wave, especially in terms of the contributions of high-order effects and non-adiabatic effects. However, it can effectively reveal the main physical characteristics of the gravitational wave signal within a single period. Especially, it is prominently manifested in describing the zoom and whirl behaviors of the periodic orbit, the influence of the hair parameter on the gravitational wave signal, and the differences from the waveform of the Schwarzschild BH. Therefore, within the scope of this study, the adiabatic approximation is reasonable and sufficient (because this paper mainly explores whether the gravitational wave signal in a single period can exhibit basic orbital characteristics such as zoom and whirl behaviors). However, the adiabatic approximation ignores the contributions of non-adiabatic effects and high-order multipole moments in the long-term orbital evolution. Specifically, non-adiabatic effects may significantly affect the amplitude and phase of the waveform on a longer time scale, especially during the long-term orbital evolution. In addition, high-order effects may also change the details of the waveform, and these effects have not been fully considered in the current adiabatic approximation.
Therefore, to depict the evolution of the gravitational wave waveform more accurately, future research will adopt more refined waveform models, especially those that can simultaneously consider high-order effects and the contributions of non-adiabatic radiation. These improvements will not only deepen our understanding of the radiation from periodic orbits but also contribute to more accurately distinguishing the properties of the short hair BH from those of the classical Schwarzschild BH. Future space-based gravitational wave detectors (such as LISA, Tianqin, Taiji, etc.) will provide crucial observational data for verifying these theoretical models. For example, by studying the gravitational wave waveform in the last year of orbital inspiral through the LISA detector and combining it with the power spectral density($S_{n}$) of the detector, it can be used to explore the influence of the hair parameter $Q_{m}$ on the long-term evolution of periodic orbits or to place constraints on it, thereby enhancing our understanding of the characteristics of the short hair BH. These issues will be further explored in future research.